OpenStudy (anonymous):

a rectangle and a triangle have equal areas. The length of the rectangle is 12 inches, and its width is 8 inches. If the base of the triangle is 32 inches, what is the length, in inches, of the altitude drawn to the base?l

OpenStudy (anonymous):

6 inches

OpenStudy (hoblos):

12*8 = 32*x/2 96 = 16x x= 6

OpenStudy (accessdenied):

\[ A_{rectangle} = b*h, A_{triangle} = \frac{1}{2}b*h\] You know the base and height of the rectangle, and the base of the triangle, so you set the areas equal and solve for the one unknown height for the triangle

OpenStudy (anonymous):

Great work guys XD